### v1

# Harnessing Test-time Adaptation for NLU tasks Involving Dialects of English

Abstract
Test-time adaptation (TTA) is an excellent method which helps generalize models across domains, tasks, and distributions without the use of labeled datasets. Thus, TTA is very useful in natural language processing (NLP) in the dialectal setting, since oftentimes, models are trained on Standard American English (SAE), evaluated on Indian English or Nigerian English, of which distribution differs significantly from the former. This is especially useful since dialectal datasets are scarce. In this paper, we explore one of the most famous TTA techniques, SHOT, in dialectal NLP. We finetune and evaluate SHOT on different combinations of dialectal GLUE. Our findings show that SHOT is a viable technique when labeled datasets are unavailable. We also theoretically propose the concept of dialectal gap and show that it has a positive correlation with the effectiveness of SHOT. We also find that in many cases, finetuning on SAE yields higher performance than finetuning on dialectal data. Our code is available at: https://github.com/dukenguyenxyz/dialect-adaptation.
Harnessing Test-time Adaptation for NLU tasks Involving Dialects of English
Duke Nguyen, Aditya Joshi, Flora Salim University of New South Wales, Australia Correspondence: research@itsduke.me
1 Introduction
Human languages exhibit natural variations with respect to four generally accepted dimensions: diaphasic (w.r.t. the medium of communication, e.g., texting versus oral communication), diastratic (w.r.t. social groups, e.g., gender), diachronic (w.r.t. time, e.g., 1960s versus 2010s English), and diatopic (w.r.t space, e.g., Indian versus Singaporean English) Zampieri et al. (2020). These variations occur primarily in three domains: orthography, syntax, and vocabulary. Examples of diatopic varieties include: "I might could help you with that“ (Southern US, Australian, and New Zealand English), "Inside tent can not see leh!“ (Singaporean English) Joshi et al. (2025). In this paper, we strictly study diatopic varieties, and we interchangeably refer to them as ‘dialects’
These natural linguistic variations pose significant challenges to NLP. Specifically, existing works have shown serious gaps with Large Language Models when processing Standard American English versus other forms of English. These gaps have been shown for bias in hate speech datasets Sap et al. (2019), dependency parsing Blodgett et al. (2018), automated speech recognition Koenecke et al. (2020), syntactic processing Ziems et al. (2023), exhibition of harmful stereotypes of speakers of AAE Hofmann et al. (2024), exhibition of stereotyping, demeaning content, etc. towards speakers of non-SAE dialects Fleisig et al. (2024).
Another challenge is the significant number of English varieties. Kortmann et al. (2020) alone have documented 77 diatopic varieties of English, however DialectBench Faisal et al. (2024), the most developed dialectal benchmark, only cover 19 varieties, many tasks of which are not equivalent. Building benchmarks and datasets also requires a significant amount of expense and annotation, which makes fine-tuning on labeled datasets more difficult to achieve.
Recent trends in computer vision (CV), however, shows a promising new field, test-time adaptation (TTA), which can overcome these challenges. Test-time adaptation can potentially adapt an LLM trained on one dialect to another unseen dialect during inference time. In this paper, we study the use of one such TTA method, SHOT Liang et al. (2020), in dialectal adaptation. We are the first to study dialectal adaptation using a TTA technique. We perform our evaluations across four dialects: SAE, Indian, Nigerian, and Singaporean across selected dialectal GLUE tasks. We compare SHOT performance against in-dialectal finetuning and cross-dialectal finetuning. Our results indicate that: (1) the wider the dialectal gap, the more effective SHOT is; (2) SHOT consistently improves the performance of models finetuned on SAE datasets and commonly outperforms models in-dialectal finetuned models; (3) cross-dialectal finetuning sometimes outcompetes in-dialectal finetuning.
2 Related Work
There are limited English dialectal benchmark datasets. DialectBench Faisal et al. (2024) is large-scale NLP dialectal benchmark covering 10 text-level tasks and 281 varieties across languages. Mutli-VALUE Ziems et al. (2023) is a dialectal translation system which applies documented syntactic structure of dialects as studied by Kortmann et al. (2020). Due to DialectBench Faisal et al. (2024)’s limited tasks in English dialects, we have decided to use Multi-VALUE Ziems et al. (2023) for our evaluation.
To overcome the dialectal gap, Dacon et al. (2022) apply adversarial learning to ensure consistence label prediction between the SAE and AAE inputs. Tan et al. (2020) propose a novel tokenization method which injects grammatical dialectical information as special symbols whilst reducing inflected words to their base form, improving the performance of AAE on SQUAD and MNLI. Held et al. (2023) use an architecture of dialectal adapters and task adapters via a contrastive loss and morphosyntactic loss. Xiao et al. (2023) use task-agnostic low-rank adapters to generalize an LLM to unseen dialects by leveraging expert linguistic knowledge via typological feature vectors. Srirag et al. (2024) similarly propose a task-agnostic low-rank adapter approach to the decoder model usecase. However, both rely on existing dialectal datasets.
We first consider domain $D$ of a joint distribution $p(x,y)$ where $x\in X,y\in Y$ are the input and output space, $p_{S}(x,y)$ and $p_{T}(x,y)$ are the distributions of the source and the target domain respectively. During test-time, we are given the source data $D_{S}={(x_{1},y_{1}),...,(x_{n}s,y_{n}s))}$ and the unlabeled target data $D_{T}={x_{1},...x_{n}t}$ Liang et al. (2023). TTA aims to use self-supervised learning on $D_{T}$ to improve the performance of the model originally trained on $D_{S}$ without access to $D_{S}$. The most famous of TTA techniques is $SHOT$ Liang et al. (2020), which leverages the use of self-supervised pseudo-labeling in conjunction with information maximization in order to improve the performance of the model at test time. Previous works have shown promise in using TTA in the NLP domain Banerjee et al. (2021); Ye et al. (2022); Zhang et al. (2021); Wang et al. (2021); Ben-David et al. (2022); Laparra et al. (2021). In addition, most have not adapted successful TTA techniques such as SHOT Liang et al. (2020), except for Zhang et al. (2021) in domain adaptation.
Task Train Eval
Indian Nigerian Singaporean Average
F TTA $\Delta$ F TTA $\Delta$ F TTA $\Delta$ F TTA $\Delta$
CoLA SAE 15.55 18.28 2.73 19.51 21.53 2.02 5.46 8.92 3.47 13.51 16.24 2.74
Indian 12.93 10.41 13.57 3.16 2.16 4.09 1.93 8.50 8.83 2.54
Nigerian 11.57 15.42 3.85 18.01 6.44 5.98 -0.47 12.01 10.70 1.69
Singaporean 13.03 15.76 2.73 13.73 12.47 -1.26 13.03 13.26 14.12 0.74
Average 13.27 16.49 3.10 15.42 15.86 1.31 6.77 6.33 1.64 11.82 12.89 2.02
SST-2 SAE 87.52 87.89 0.37 88.01 88.89 0.88 86.39 86.64 0.25 87.31 87.81 0.50
Indian 89.13 88.64 90.14 1.50 85.64 87.77 2.13 87.80 88.96 1.82
Nigerian 86.02 87.02 1.00 89.63 85.39 85.77 0.38 87.01 86.40 0.69
Singaporean 87.52 86.64 -0.88 87.52 88.26 0.74 88.38 87.81 87.45 -0.07
Average 87.55 87.18 0.16 88.45 89.10 1.04 86.45 86.73 0.92 87.48 87.67 0.71
RTE SAE 58.57 59.76 1.19 58.96 60.16 1.20 58.17 58.96 0.79 58.57 59.63 1.06
Indian 54.98 54.18 52.99 -1.19 53.78 54.98 1.20 54.31 53.99
Nigerian 58.96 55.78 -3.18 58.96 57.77 56.57 -1.20 58.56 56.18 -2.19
Singaporean 54.98 54.98 60.16 61.75 1.59 60.16 58.43 58.37 0.80
Average 56.87 56.84 -0.66 58.07 58.30 0.53 57.47 56.84 0.26 57.47 57.04 -0.08
Table 1: Matrix comparison of models fine-tuned and applied with TTA on different dialects then evaluated on other dialects. F stands for fine-tuning, $\Delta=TTA-F$ is the performance difference before and after TTA is applied on a model that has been fine-tuned on the evaluation dialect. Bold is for best value for specific task and setting. CoLA SAE on SAE performance is 49.44. SST-2 SAE on SAE performance is 91.97. RTE SAE on SAE performance is 58.12.
3 Method
Our model architecture is adapted from Liang et al. (2020) for NLP [^1] Our SHOT implementation is based on Liang et al. (2020) at https://github.com/tim-learn/SHOT, taking inspiration from Zhang et al. (2021). Specifically, inputs are first embedded via a frozen transformer-based Encoder-only model $F$, followed by a bottleneck network $B$, and then a classifier layer $C$. We choose our $F$ as roberta-base Liu et al. (2019), $B$ as a 2-layer bi-directional GRU Cho et al. (2014), and $C$ being 2-layer fully connected. Our optimizer is AdamW Loshchilov and Hutter (2019) with annealing strategy for the learning rate Long et al. (2017) adjusted during propagation with the formula $\eta_{p}=\frac{\eta_{0}}{(1+10p)^{0.75}}$, $\eta_{0}=0.001$, $p$ being the training progress linearly increasing from 00 to $1$. Our architecture is based on Zhang et al. (2021).
Dialectal gap is the distribution shift between two dialects. There has been extensive work in concept drift and distribution shift literature. However, methods to measure distribution shift in textual datasets have not been explored previously. To analyze distribution shift between dialects, we propose dialectal gap, measured as the difference between the evaluative performance on a dialect A and a set of dialects {B}. Our formulation of the dialectal gap is a modification over Faisal et al. (2024). Existing literature Srirag et al. (2024) shows that popular LLMs are not reliable models to measure the best optimal performance of dialects due to their pretraining data being predominantly SAE. Hence, we will only quantify the dialectal gap with A being SAE. Given that $eval(X,Y)\in[0,1]$ outputs the performance of a model trained on dialect $X$ evaluating on dialect $Y$ We can thus measure the dialectal gap as $eval(A,A)-eval(A,B_{i})$ and the average dialectal gap as:
$\frac{1}{n}\sum^{n}_{i=1}eval(A,A)-eval(A,B_{i})$ (1)
We hypothesize that there is a relationship between the dialectal gap and $\Delta$. Hence we perform Pearson correlation analysis between these two variables. We also perform statistic comparisons to analyze the relationship between dialect via finetuning and TTA in Section 4.
4 Experiments
Eval
Indian Nigerian Singaporean Average
Train B A $\Delta$ B A $\Delta$ B A $\Delta$ B A $\Delta$
SAE 67.96 68.93 0.98 68.91 69.94 1.03 65.76 66.69 0.92 67.54 68.52 0.98
Indian 66.86 66.01 66.64 0.63 63.50 64.93 1.43 65.46 65.79 1.03
Nigerian 66.92 66.84 -0.09 69.20 65.46 65.11 -0.35 67.19 65.97 -0.22
Singlish 66.34 66.50 0.16 68.18 68.75 0.57 68.35 67.62 67.62 0.36
Average 67.02 67.42 0.37 68.08 68.44 0.60 65.77 65.58 0.32 66.95 67.15 0.43
Table 2: Average evaluation on CoLA, RTE, and SST-2. Bold is for best value for specific task and setting. Italic is for best performing evaluative dialect per dialect trained.
Dialect-finetuned SAE-finetuned SHOT on SAE SAE-finetuned - SHOT Dialect-finetuned - SHOT
CoLA SST-2 RTE CoLA SST-2 RTE CoLA SST-2 RTE CoLA SST-2 RTE CoLA SST-2 RTE
Indian 12.93 89.13 54.98 15.55 87.52 58.57 18.28 87.89 59.76 2.73 0.37 1.19 -5.35 1.24 -4.78
Nigerian 18.01 89.63 58.96 19.51 88.01 58.96 21.53 88.89 60.16 2.02 0.88 1.20 -3.52 0.74 -1.20
Singaporean 13.03 88.38 60.16 5.46 86.39 58.17 8.92 86.64 58.96 3.47 0.25 0.79 4.11 1.74 1.20
Average 14.66 89.05 58.03 13.51 87.31 58.57 16.24 87.81 59.63 2.74 0.50 1.06 -1.59 1.24 -1.59
Table 3: Comparison of Finetuning on Dialect (Dialect-finetuned) versus Finetuning on SAE (SAE-Finetuned) versus SHOT. Difference between SAE-Finetuned and SHOT compared against difference betwen Dialect-Finetuned and SHOT.
We perform our experiments on selected GLUE Wang et al. (2019) tasks: CoLA Warstadt et al. (2019), RTE Wang et al. (2019), SST-2 Socher et al. (2013) with dialectal transformation via Multi-VALUE Ziems et al. (2023). The dialects which we select are based on a subset from Ziems et al. (2023): SAE (Standard American English), Indian (Indian English), Nigerian (Nigerian English), and Singaporean (Singaporean English).
Our implementation of SHOT Liang et al. (2020) follows the default hyperparameters, and otherwise Zhang et al. (2021). Each training was done on 6 NVIDIA RTX A5000. All of our hyperparameters are shown in Table 5. Experimental results are shown in Table 1. $\Delta$ stands for the performance difference before and after TTA is applied on a model that has been fine-tuned on the evaluation dialect. We calculate the individual dialectal gap according to Equation 1 in Table 4.
We perform Pearson correlation analysis of Gap against $\Delta$, we obtain a p-value of $0.0041<0.005$, indicating that there is a strong relationship between the two and the coefficient is $0.8455$, suggesting that there is a positive relationship between gap and $\Delta$. Based on this evidence, we observe that there is a strong positive relationship between the dialectal gap and $\Delta$. This means that as the dialectal gap increases, the more TTA will be beneficial in reducing said gap.
Indian Nigerian Singaporean
CoLA Gap 16.95 14.97 21.99
$\Delta$ 2.73 2.02 3.47
SST-2 Gap 4.45 3.96 5.58
$\Delta$ 0.37 0.88 0.25
RTE Gap -0.45 -0.84 -0.05
$\Delta$ 1.19 1.20 0.79
Table 4: Dialectal gap. CoLA has a range of $[-1,1]$ so we rescaled it to $[0,1]$. All values are in percentage. Dialectal gaps are calculated with respect to SAE. $\Delta$ is between the TTA and the baseline model as finetuned on SAE and evaluated on the corresponding dialect.
We also compare the performance of using SHOT versus performing finetuning on dialectal data. Table 3 shows that SHOT always improves the performance of models trained on SAE data. Furthermore, SHOT commonly out-performances model trained on dialectal data. This suggests that building dialect-robust models require the incorporation of SHOT-based self-supervised techniques in order to reverse the effect of performing self-supervised learning of base embedding models on SAE data.
When analyzing the relationship between dialects models are finetuned on and their performance on varying dialects, we also observe the asymmetry of evaluation. Specifically, as shown on Table 2, we find that when training on a dialect $A$, its performance is best when evaluating on dialect $A$ (with the exception of SAE, see italicised entries on Table 2). However, when training on dialect $A$, $B_{1}$, $B_{2}$, etc., and evaluating on dialect $A$, oftentimes, training on a dialect $B_{i}$ yields higher performance than training on dialect $A$. Specifically, SAE shows the best performance with or without test-time adaptation in multiple dialects (see boldened entries in Table 2).
5 Conclusion
Due to the dialectal gap which exhibits in previous literature, we formally propose a mathematical definition of the concept, apply SHOT Liang et al. (2020) to dialectal NLP, and do a comparative study of SHOT against finetuning under different dialectal settings. Our findings indicate that: (1) There is a strong positive relationship between the dialectal gap and $\Delta$. This means that as the dialectal gap increases, the more TTA will be beneficial in reducing said gap. (2) SHOT always improves the performance of models trained on SAE data. Furthermore, SHOT commonly out-performances model trained on dialectal data. (3) When training on a dialect $A$, its performance is best when evaluating on dialect $A$. However, when training on dialect $A$, $B_{1}$, $B_{2}$, etc., and evaluating on dialect $A$, oftentimes, training on a dialect $B_{i}$ yields higher performance than training on dialect $A$.
Limitations
We are unable to perform evaluation on other GLUE tasks due to their dataset sizes being significantly higher than our computational requirement. We also weren’t able to do experiments on STS-B due to this being a regression task which is not yet accommodated by SHOT Liang et al. (2020). We use Multi-VALUE to create dialectal transformations. We acknowledge that there may be errors. Similarly, dialects are a combination of syntax, vocabulary, orthography and culture. These nuances are not captured in our evaluative datasets. RoBERTA was pretrained predominantly on SAE corpus, we have not explored the effect of RoBERTA or other language models if pretrained with dialectal data predominantly.
Acknowledgments
The production of the code used in this paper was partially reliant on generative AI code assistant. Generative AI tools were also used to perform literature search. All AI-generation has been validated by the authors. This work was funded by Duke Nguyen’s PhD Scholarship at the University of New South Wales, Australia. Dedicated to the memory of Ryan Herford (1997-2024).
Appendix A Appendix
Hyperparameters
Epochs 30
Batch size (per device) 32
Label smoothing 0.1
IM loss multiplier 1.0
Classifier loss multiplier 0.3
Entropy epislon 1e-5
Table 5: Hyperparameters for finetuning and SHOT